# Decoding Diffusion: A Scalable Framework for Unsupervised Analysis of Latent Space Biases and Representations Using Natural Language Prompts

### v1

@@ 4.3 Cluster Visualization
Figure 4: t-SNE visualization of h-space vectors sampled from the Food500-CAP dataset. Clusters are labelled based on the id assigned by HDBSCAN [30]. Some prominent clusters include (8) square dishes; (12) pots with soup; (27) sandwich; (44) pie; (94) takeout boxes; (98) rectangular plates; (120) salads; (127) stir fried noodles; and (144) seafood boil; (136) purple vegetables, esp. purple cabbage
@@
Figure 5: (a) default image for “a photo of food” with seed 0, (b) image conditioned on “a photo of food” plus the average of cluster 1 from Figure 4, depicting a square plate (c) image conditioned on “a photo of food” plus the average of cluster 2 from Figure 4, depicting bread (d) image conditioned on “a photo of food” plus the average of cluster 3 from Figure 4, depicting a pots (e) image conditioned on “a photo of food” plus the average of clusters 1 and 3 from Figure 4, depicting a square pot

### v2

@@ 4.3 Cluster Visualization
Figure 4: t-SNE visualization of h-space vectors sampled from the Food500-CAP dataset. Clusters are labelled based on the id assigned by HDBSCAN [30]. Some prominent clusters include (8) square dishes; (12) pots with soup; (27) sandwiches/bread; (44) pie; (94) takeout boxes; (98) rectangular plates; (120) salads; (127) stir fried noodles; and (144) seafood boil; (136) purple vegetables, esp. purple cabbage
@@
Figure 5: (a) default image for “a photo of food” with seed 0, (b) image conditioned on “a photo of food” plus the average of cluster 8 from Figure 4, depicting a square plate (c) image conditioned on “a photo of food” plus the average of cluster 27 from Figure 4, depicting bread (d) image conditioned on “a photo of food” plus the average of cluster 12 from Figure 4, depicting a pots (e) image conditioned on “a photo of food” plus the average of clusters 8 and 12 from Figure 4, depicting a square pot